OpenStudy (anonymous):

Match the sequence with the correct indicated term a12 for 8,3,-2,...

OpenStudy (anonymous):

Choose... -175 340 -47 -25/2 25/2

OpenStudy (anonymous):

again an AP \[a _{n}=a+(n-1)d \implies a _{12}=8+(12-1)(-5)\implies a _{12}=8+(11)(-5)\]

OpenStudy (anonymous):

\[a _{12}=8-55=-47\]